Solve the second order differential equation for an LRC circuit in series with the initial conditions given. And create a graph

(100x10->) + (470) ** (470x10-69 Con las condiciones iniciales siguientes: g(0)=0 y (0)=0.
Answer #1

(100 x103) q15) + 470 dit) + 1 9 =9 (470x1062 =9 0.1 7 (+) + 470 aCk) +21.27-669=9 The ausiliary equation is : 0.1 m² + fro910) =0>):6, ++4230x17?=0 ».64 + 6 =-4-2308103- 910).=0 =) -4695.450,-4-555,=0 0x4695.45 -> 4695-456+4695-45; = ...:-19-86 –4